Reports 1-1 of 1 Clear search Modify search
AdV-TCS (TCS control system)
mantovani, nardecchia, degallaix, tacca, rocchi - 21:43 Friday 20 May 2022 (55923) Print this report
TCS (CH+DQS) tuning on the North arm

Summary of the work on the North arm, similar to the one on the West arm the day before, see post 55908.

The summary of the day is in the plot attached:

- at the start switch on the CH from 0 to 160 minutes
- then fine tuning of the DAS to restore the matching from 160 minutes to 520 minutes where different setting are tried
- then at 520 minutes switch off the inner DAS to create a big transient to scan fast the CP length (sucessfull to see the change in the HWS)
- based on those data, 630 minutes new setting of the DAS but still not enough recover the matching.

The script to analyse the data have been updated and are in the daily folder


Images attached to this report
Comments to this report:
rocchi - 22:02 Friday 20 May 2022 (55925) Print this report

Here is also the HWS-DET curvature data with the indication of all the steps done today. The figure shows that the best matching is reached when the curvature is closer to zero. The spike before each DAS step is due to the flipper mirror being closed to perform the power tuning.

We leave the NI DAS with the current settings: 1.9 W inner ring; 4.2 W outer ring.

Images attached to this comment
Search Help